Find the equation of a line containing points (2,6) and (8,9).
Answer by ewatrrr(24785)


Hi,
Using the slope-intercept general form of an equation of a line:
y = mx + b where m is th eslope and b the y-intercept
points (2,6) and (8,9) on the line
slope is = (9-6)/(8-2) = 3/6 = 1/2
y = (1/2)x + b Using (2,6) to solve for b
6 = 1 + b
5 = b
y = (1/2)x + 5 the equation of the line